Calculate required rate of return “r and Calculate intrinsic value. [Assume or substitute Net Income for Cash Flow].

Based on the following information:

Summary Income Statement in $

Sales                5,000,000

EBIT               1,000,000

Int. Expense        40,000

EBT                   960,000

Net Income        624,000

Other Financial Information in $

Market Value of LT Debt          500,000

Market Value of Equity          4,500,000

Avg. P/E                               12x

Growth rate “g” = 0.04
Answer #1
1 Growth Rate = 0.04 or 4 %
Growth Rate = ROE x Retention Ratio = 4 %
(Net Income / Market Value of Equity) x 100 x Retention Ratio = 4 %
(624000 / 4500000) x 100 x Retention Ratio = 4
13.867 x Retention Ratio = 4
Retention Ratio = 4 /13.867 ~ 0.2884
Dividend Payout Ratio = (1-Retention Ratio) = (1-0.2884) = 0.7116
Therefore, Dividend Paid = Net Income x Dividend Payout Ratio = 624000 x 0.7116 = $ 444038.4
Expected Future Dividend = (1+Growth Rate) x Dividend Paid = (1.04) x 444038.4 = $ 461799.936
Expected Dividend per Share = D1 = 461799.936 / K (K is the assumed total number of shares outstanding) and Current Price Per Share = P0 = Market Value of Equity / K = 4500000 / K
Therefore, Required Rate of Return = (D1/P0) + g = (461799.936/K ) x (K / 4500000) + 0.04 = 0.10262 + 0.04 = 0.14262 or 14.262 %
2 Net Income = Cash Flow (assumed) = $ 624000 and perpetual growth rate = 4 %
Required Rate of Return = 14.262 %
Expected Cash Flow = 624000 x 1.04 = $ 648960
Therefore, Intrinsic Value = 648960 / (0.14262 - 0.04) = $ 6323777.40